OpenStudy (anonymous):

Rewrite 5/ 16 + 6/ 20 with a common denominator.

OpenStudy (anonymous):

\[\frac{ 5 }{ 16 } + \frac{ 6 }{ 20 } =\frac{ 25+24 }{ 80 }=49/80\]